Find the sum of 3/4+3/28+3/70+3/130+ upto 3/9700 can you explain me how to solve this and what are it's steps followed and it's formulae

Grade:7

1 Answers

Arun
5 years ago
 
3/4=1-1/4
3/28=1/4-1/7
3/70=1/7-1/0
…..
3/9700 = 1/97-1/100
Add all the above..What remains is 1-1/100 = .99
 
Regards
Arun (askIITians forum expert)